# Connected components in d-minimal structures

### v1

@@ 1. Introduction
Every connected component of a set definable in an o-minimal structure over $\mathbb{R}$ is definable, but this is not the case for locally o-minimal structures and d-minimal structures over $\mathbb{R}$. For instance, [1, Theorem 1.9] gives a necessary and sufficient condition for every connected component of a definable set to be definable for locally o-minimal expansions of $(\mathbb{R},<,+,\mathbb{Z})$. $(\mathbb{R},<,+,\mathbb{Z})$ has a definable set whose connected component is not definable. The structure $(\mathbb{R},+,\cdot,2^{\mathbb{Z}})$ is a d-minimal structure by [2], but there exists a set definable in this structure one of whose connected component is not definable [7, Example 79]. A question is whether a given tame structured $\mathfrak{R}$ can be expanded to a structure $\mathfrak{R}^{\natural}$ so that every connected component of a set definable in $\mathfrak{R}$ is definable in $\mathfrak{R}^{\natural}$ without loosing the given tameness such as local o-minimality and d-minimality.
@@ 2. Multi-cell decomposition in $\mathfrak{R}$
We prove a multi-cell decomposition theorem, which refines partition of definable sets into special submanifolds in d-minimal structures [5, 8]. This theorem is used in the proof of a generalized version of Theorem 1.1. Throughout this section, suppose that $\mathfrak{R}=(\mathbb{R},<,+,\ldots)$ is a d-minimal expansion of the ordered group of reals. ‘Definable’ means ‘definable in $\mathfrak{R}$ with parameters’. For a subset $S$ of $\mathbb{R}^{n}$, $\operatorname{int}(S)$, $\operatorname{cl}(S)$ and $\operatorname{fr}(S):=\operatorname{cl}(S)\setminus S$ denote the interior, closure and frontier of $S$, respectively. Let $\pi:\mathbb{R}^{n}\to\mathbb{R}^{n-1}$ be the coordinate projection forgetting the last coordinate. For a subset $S$ of $\mathbb{R}^{n}$, $\pi|_{S}$ denotes the restriction of $\pi$ to $S$.
@@
    $Y$ =$=\{(x,y)\in\pi(Y)\times\mathbb{R}\;|\;y=f(x)\}\text{,}$
    $Y$ =$=\{(x,y)\in\pi(Y)\times\mathbb{R}\;|\;y>f(x)\}\text{,}$
    $Y$ =$=\{(x,y)\in\pi(Y)\times\mathbb{R}\;|\;y<g(x)\}\text{ and }$
    $Y$ =$=\{(x,y)\in\pi(Y)\times\mathbb{R}\;|\;f(x)<y<g(x)\}$
@@
Observe that every bounded pseudo-special submanifolds enjoys property ()($\dagger$)( † ).
@@
    For every $n\in\mathbb{N}$ and every pseudo-special submanifold $M$ of $\mathbb{R}^{n}$, property ()($\dagger$)( † ) is fulfilled.
@@
Let $C_{1},C_{2}\subseteq\mathbb{R}^{n}$ be two connected graphs on $V$ with $\underline{C_{1}}<_{V}\underline{C_{2}}$. We denote $\{(x,t)\in V\times\mathbb{R}\;|\;f_{C_{1}}(x)<t<f_{C_{2}}(x)\}$ by $(C_{1},C_{2})_{V}$ and $\{(x,t)\in V\times\mathbb{R}\;|\;f_{C_{1}}(x)\leq t\leq f_{C_{2}}(x)\}$ by $[C_{1},C_{2}]_{V}$.
@@
A well-ordered pair $(M,P)$ is an extremely well-ordered pair if $\operatorname{Adh}^{+}(M,P)\cap\pi^{-1}(x)=\operatorname{Adh}^{+}(M,P,x)$ and $\operatorname{Adh}^{+}(M,P)\cap\pi^{-1}(x)=\operatorname{Adh}^{+}(M,P,x)$ for every $x\in\pi(M)$.
@@
$\quad((B\cap P\text{ is the graph of a continuous function }f\text{ defined on }\pi(P)\cap\pi(B))$
@@
  • $\operatorname{iso}_{m}(S)=\operatorname{iso}(\operatorname{iso}_{m-1}(S))$;
@@
    $M$ enjoys property ()($\dagger$)( † ).
@@
    If $N>1$, for every $1\leq i<N$, the definable set
    $P_{i}=\bigcup_{x\in\pi(M)}\operatorname{iso}_{i}(\operatorname{cl}(M)\cap\pi^{-1}(x))$
    is a multi-cell. Observe that this condition implies that $(M,P_{i})$ and $(P_{j},P_{k})$ are well-ordered pairs for $11\leq i\leq N-$ 1 and $11\leq j<k\leq N-$ 1;
@@
    $(M,P_{i})$ and $(P_{j},P_{k})$ are extremely well-ordered pairs for $11\leq i\leq N-$ 1 and $11\leq j<k\leq N-$ 1.
@@
(a) First we show uniqueness. Let $\beta_{1},\beta_{2}:[0,1]\to X$ be continuous maps with $\alpha=\pi\circ\beta_{i}$ and $\beta_{i}(0)=b$ for $i=1,2$. Since $M$ is a pseudo-special submanifold, $I:=\{t\in[0,1]\;|\;\beta_{1}(t)=\beta_{2}(t)\}$ is a closed and open subset of $[0,1]$ containing 00. We have $I=[0,1]$ because $[0,1]$ is connected.
Consider the set $J$ of $t\in[0,1]$ such that there exists a continuous map $\beta_{t}:[0,t]\to X$ with $\beta_{t}(0)=a$ and $\pi\circ\beta_{t}=\alpha|_{[0,t]}$. $J$ is open because $M$ is pseudo-special. Let $t_{0}:=\sup J$. For every $0\leq t<t_{0}$, there exists a unique continuous map $\beta_{t}:[0,t]\to X$ satisfying the above conditions. The map $\beta:[0,t_{0})\to X$ defined by $\beta(t):=\beta_{t}(t)$ is a continuous map by uniqueness of $\beta_{t}$. Let $(s_{m})_{m\in\mathbb{N}}$ be an increasing sequence in $[0,t_{0})$ converging to $t_{0}$. By property ()($\dagger$)( † ), we may assume that $(\beta(s_{m}))_{m\in\mathbb{N}}$ is a convergent sequence by taking a subsequence. Put $c=\lim_{m\to\infty}\beta(s_{m})$. We have $\pi(c)=\alpha(t_{0})\in\pi(M)$. Therefore, we obtain $c\in\operatorname{cl}(X)\cap\pi^{-1}(\pi(M))$.
@@
We may assume that condition (i) is satisfied without loss of generality. By taking $u_{0}<t_{0}$ sufficiently close to $t_{0}$, we may assume that $\alpha([u_{0},t_{0}])$ is contained in $V^{\prime}$. By taking a subsequence of $(s_{m})_{m\in\mathbb{N}}$, we may assume that $\beta(s_{m})\in X\cap(Z^{\prime},C_{1})_{V^{\prime}}$ and $u_{0}<s_{m}<t_{0}$ for every $m\in\mathbb{N}$. By using the assumption that $M$ is a pseudo-special submanifold and condition (i), we can easily prove that $K:=\{t\in[s_{1},t_{0})\;|\;\beta(t)\in X\cap(Z^{\prime},C_{1})_{V^{\prime}}\}$ is closed and open in $[s_{1},t_{0})$. We omit the details. Since $s_{1}\in K$, we have $K=[s_{1},t_{0})$. This means that $\beta([s_{1},t_{0}))\subseteq X\cap(Z^{\prime},C_{1})_{V^{\prime}}$. Since $\beta([s_{1},t_{0}))$ is connected, it is contained in a connected component of $X\cap(Z^{\prime},C_{1})_{V^{\prime}}$, say $D$. However, $D$ is a connected graph on $V^{\prime}$ with $Z^{\prime}<_{V^{\prime}}D$ by condition (i). We have $c\in\operatorname{cl}(D)$, but $c\in Z^{\prime}$ and $Z^{\prime}\cap\operatorname{cl}(D^{\prime})=\emptyset$, which is absurd.
@@
Assume for contradiction that $s_{0}:=\sup K<1$. Since $M$ is pseudo-special and uniqueness of $\widetilde{H}$, for every $0\leq t\leq 1$, there exist $\delta_{t}>s_{0}$ and $0\leq\varepsilon_{1,t}<t<\varepsilon_{2,t}\leq 1$ such that the restriction of $\widetilde{H}$ to $[0,\delta_{t}]\times[\varepsilon_{1,t},\varepsilon_{2,t}]$ is continuous. Since $[0,1]$ is compact, there exists $\delta>s_{0}$ such that the restriction of $\widetilde{H}$ to $[0,\delta]\times[0,1]$ is continuous, which is absurd. ∎
We show $\pi(X)=V$. Since $\pi|_{X}$ is a local homeomorphism, $\pi(X)$ is open in $V$. Assume for contradiction that $\pi(X)$ is not closed in $V$. We can find $a\in\operatorname{fr}(\pi(X))\cap V$. Let $(a_{m})_{m\in\mathbb{N}}$ be a convergent sequence in $\pi(X)$ converging to $a$. Take $b_{m}\in X$ with $\pi(b_{m})=a_{m}$. Since $M$ enjoys property ()($\dagger$)( † ), we may assume that $(b_{m})_{m\in\mathbb{N}}$ is convergent in $\mathbb{R}^{n}$ by taking a subsequence. Let $b:=\lim_{m\to\infty}b_{m}$. We have $\pi(b)=a\notin\pi(X)$ and $b\notin X$. This implies that $b\in\operatorname{fr}X\cap\pi^{-1}(\pi(M))$ and $\pi(b)\notin\pi(X)$. These contradict the either Claim 1 or ‘in particular’ part of Claim 2.
@@
[…]
We show this theorem by induction on $n$. Suppose $n=1$. Let $J=\operatorname{int}(M)$. Let $N=\operatorname{rank}(\operatorname{cl}(M\setminus J))$ and set $S_{i}:=\operatorname{iso}_{i}(\operatorname{cl}(M\setminus J))\cap(M\setminus J)$ for $10\leq i\leq N-$ 1. It is easy to show that $\{J,S_{0},\ldots,S_{N-1}\}$ is a multi-cell decomposition of $M$ after removing empty sets in the collection. We omit the details.
@@
Put $M[i]:=M\cap\pi^{-1}(\{x\in\pi(M)\;|\;\operatorname{rank}(\operatorname{cl}(M\cap\pi^{-1}(x)))=i\})$ for $1\leq i\leq N$. We may assume that $\operatorname{rank}(\operatorname{cl}(M\cap\pi^{-1}(x)))$ is independent of $x\in\pi(M)$, say $N$, by considering $M[i]$ instead of $M$. We define $P_{k}$ for $11\leq k\leq N-$ 1 as in clause (5) of Lemma 2.4. Put $P_{0}:=M$ for simplicity of notation. We prove the following claim which implies a multi-cell decomposition partitioning $M$.
Claim 1. There exists a multi-cell decomposition $\{C_{1},\ldots,C_{l}\}$ partitioning $\pi(M)$ such that, for $1\leq i\leq l$, $P_{i}\cap\pi^{-1}(C_{j})$ is a multi-cell for $10\leq i\leq N-$ 1 and clause (6) in Lemma 2.4 is satisfied for $M\cap\pi^{-1}(C_{i})$.
@@
Suppose $N>1$. Apply the induction hypothesis to $P_{1}$, then we may assume that $P_{1},\ldots,P_{N-1}$ are multi-cells and $(P_{i},P_{j})$ are extremely well-ordered pairs for $11\leq i<j\leq N-$ 1. Observe that $\pi(M)$ is a multi-cell because $\pi(M)=\pi(P_{1})$ and $P_{1}$ is a multi-cell. We want to reduce to the case where $(M,P_{i})$ is a well-ordered pair for every $11\leq i\leq N-$ 1. Put $T_{1i}:=\{x\in\pi(M)\;|\;\operatorname{Adh}^{+}(M,P_{i})\cap\pi^{-1}(x)\neq\operatorname{Adh}^{+}(M,P_{i},x)\}$ and $T_{2i}:=\{x\in\pi(M)\;|\;\operatorname{Adh}^{+}(M,P_{i})\cap\pi^{-1}(x)\neq\operatorname{Adh}^{+}(M,P_{i},x)\}$. Put $T:=\bigcup_{i=1}^{N-1}T_{1i}\cup T_{2i}$. $T$ is nowhere dense by Lemma 2.3.
@@
A multi-cell decomposition partitioning $\{V_{+\infty},V_{+\infty},V_{\mathbb{R}},W\}$ is a multi-cell decomposition partitioning $M$. Therefore, we have only to construct multi-cell decompositions partitioning each of $V_{+\infty}$, $V_{+\infty}$, $V_{\mathbb{R}}$ and $W$. Observe that $V_{\mathbb{R}}=\pi(V_{\mathbb{R}})\times\mathbb{R}$. It is easy to show that $V_{\mathbb{R}}$ is a union of multi-cells. We omit the details. We next show that there exists a multi-cell decomposition partitioning $V_{+\infty}$. By considering a multi-cell decomposition of $\pi(V_{+\infty})$, we may reduce to the case where $\pi(V_{+\infty})$ is a multi-cell. Consider the definable function $f:\pi(V_{+\infty})\to\mathbb{R}$ given by $f(x)=\inf\{t\in\mathbb{R}\;|\;(x,t)\in V_{+\infty}\}$. Let $D$ be the set of points at which $f$ is discontinuous. By [5, 8.3 Almost continuity], we have $\dim D<\dim\pi(V_{+\infty})$. We can reduce to the case in which $f$ is continuous by induction on $\dim\pi(V_{+\infty})$. $V_{+\infty}$ is a multi-cell because $\pi(V_{+\infty})$ is a multi-cell and $V_{+\infty}=\{(x,t)\in\pi(V_{+\infty})\times\mathbb{R}\;|\;t>f(x)\}$. We can prove that there exists a multi-cell decomposition partitioning $V_{-\infty}$ in the same manner.
@@
Let $Q:=\{(x,y)\in\pi(W)\times\mathbb{R}\;|\;\exists z\ (x,y,z)\in Z\}$. Observe that $\pi(W)=\pi(Q)$ and $\dim Q\cap\pi^{-1}(x)=0$ for every $x\in\pi(Q)$. By Claim 1, there exists a multi-cell decomposition $\{Q_{1},\ldots,Q_{l}\}$ of $Q$. Observe that, for every $(x,y)\in Q$, there exists a unique $z\in\mathbb{R}$ with $(x,y,z)\in Z$. Let $g:Q\to\mathbb{R}$ be the definable function sending $(x,y)\in Q$ to the unique point $z\in\mathbb{R}$ with $(x,y,z)\in Z$. Put $g_{i}:=g|_{Q_{i}}$ for $1\leq i\leq l$. By [5, 8.3 Almost continuity], the set $G_{i}$ of points at which $g_{i}$ is discontinuous is nowhere dense in $Q_{i}$. For every $C\in\operatorname{Con}(Q_{i})$, $C$ is the graph of a continuous function defined on a connected component of $\pi(Q_{i})$. Therefore, $\pi(G_{i}\cap C)$ is nowhere dense in $\pi(Q_{i})$ for every $C\in\operatorname{Con}(Q_{i})$. Since $\operatorname{Con}(Q_{i})$ is a countable set, $\pi(G_{i})=\bigcup_{C\in\operatorname{Con}(Q_{i})}\pi(G_{i}\cap C)$ has an empty interior in $\pi(Q_{i})$ by the Baire category theorem. Since $\pi(G_{i})$ is definable in the d-minimal structure $\mathfrak{R}$, $\dim\pi(G_{i})<\dim\pi(Q_{i})$. By induction on $\dim(\pi(Q_{i}))$, using the induction hypothesis, there exists a multi-cell decomposition $D_{1},\ldots,D_{m}$ of $\pi(Q)$ such that $Q\cap\pi^{-1}(D_{i})$ is a multi-cell and the restriction of $g$ to $Q\cap\pi^{-1}(D_{i})$ is continuous. Since $Q\cap\pi^{-1}(D_{i})$ is a multi-cell, every connected component of $W\cap\pi^{-1}(D_{i})$ is of the form $\{(x,y)\in D_{i}\times\mathbb{R}\;|\;f(x)<y<g(x,f(x))\}$, where $f$ is a continuous function defined on $D_{i}$. This means that $W\cap\pi^{-1}(D_{i})$ is a multi-cell. ∎
@@
Let $\mathfrak{R}$ be a d-minimal expansion $(\mathbb{R},<,+,\ldots)$ of the ordered group of reals which satisfies one of the following conditions:

### v2

@@ 1. Introduction
Every connected component of a set definable in an o-minimal structure over $\mathbb{R}$ is definable, but this is not the case for locally o-minimal structures and d-minimal structures over $\mathbb{R}$. For instance, [1, Theorem 1.9] gives a necessary and sufficient condition for every connected component of a definable set to be definable for locally o-minimal expansions of (,<,+,)($\mathbb{R}$$\mathbb{Z}$). (,<,+,)($\mathbb{R}$$\mathbb{Z}$) has a definable set whose connected component is not definable. The structure $(\mathbb{R},+,\cdot,2^{\mathbb{Z}})$ is a d-minimal structure by [2], but there exists a set definable in this structure one of whose connected component is not definable [7, Example 79]. A question is whether a given tame structured $\mathfrak{R}$ can be expanded to a structure $\mathfrak{R}^{\natural}$ so that every connected component of a set definable in $\mathfrak{R}$ is definable in $\mathfrak{R}^{\natural}$ without loosing the given tameness such as local o-minimality and d-minimality.
@@ 2. Multi-cell decomposition in $\mathfrak{R}$
[…]
@@
    $Y$ $=\{(x,y)\in\pi(Y)\times\mathbb{R}\;|\;y=f(x)\}\text{,}$
    $Y$ $=\{(x,y)\in\pi(Y)\times\mathbb{R}\;|\;y>f(x)\}\text{,}$
    $Y$ $=\{(x,y)\in\pi(Y)\times\mathbb{R}\;|\;y<g(x)\}\text{ and }$
    $Y$ $=\{(x,y)\in\pi(Y)\times\mathbb{R}\;|\;f(x)<y<g(x)\}$
@@
Observe that every bounded pseudo-special submanifolds enjoys property $(\dagger)$.
@@
    For every $n\in\mathbb{N}$ and every pseudo-special submanifold $M$ of $\mathbb{R}^{n}$, property $(\dagger)$ is fulfilled.
@@
Let $C_{1},C_{2}\subseteq\mathbb{R}^{n}$ be two connected graphs on $V$ with $\underline{C_{1}}<_{V}\underline{C_{2}}$. We denote $\{(x,t)\in V\times\mathbb{R}\;|\;f_{C_{1}}(x)<t<f_{C_{2}}(x)\}$ by $(C_{1},C_{2})_{V}$ and $\{(x,t)\in V\times\mathbb{R}\;|\;f_{C_{1}}(x)\leq t\leq f_{C_{2}}(x)\}$ by $C_{1},C_{2}]_{V}$.
@@
A well-ordered pair $(M,P)$ is an extremely well-ordered pair if $\operatorname{Adh}^{+}(M,P)\cap\pi^{-1}(x)=\operatorname{Adh}^{+}(M,P,x)$ and $\operatorname{Adh}^{-}(M,P)\cap\pi^{-1}(x)=\operatorname{Adh}^{-}(M,P,x)$ for every $x\in\pi(M)$.
@@
[…]
@@
  • $\operatorname{iso}_{m}(S)=\operatorname{iso}(S\setminus\bigcup_{i=0}^{m-1}\operatorname{iso}_{i}(S))$;
@@
    $M$ enjoys property $(\dagger)$.
@@
    If $N>1$, for every $1\leq i<N$, the definable set P_i=_x π(M) iso_i(cl(M) π^-1(x)) is a multi-cell. Observe that this condition implies that $(M,P_{i})$ and $(P_{j},P_{k})$ are well-ordered pairs for $1\leq i\leq N-1$ and $1\leq j<k\leq N-1$;
@@
    $(M,P_{i})$ and $(P_{j},P_{k})$ are extremely well-ordered pairs for $1\leq i\leq N-1$ and $1\leq j<k\leq N-1$.
@@
(a) First we show uniqueness. Let $\beta_{1},\beta_{2}:[0,1]\to X$ be continuous maps with $\alpha=\pi\circ\beta_{i}$ and $\beta_{i}(0)=b$ for $i=1,2$. Since $M$ is a pseudo-special submanifold, $I:=\{t\in[0,1]\;|\;\beta_{1}(t)=\beta_{2}(t)\}$ is a closed and open subset of $[0,1]$ containing 0. We have $I=[0,1]$ because $[0,1]$ is connected.
Consider the set $J$ of $t\in[0,1]$ such that there exists a continuous map $\beta_{t}:[0,t]\to X$ with $\beta_{t}(0)=a$ and $\pi\circ\beta_{t}=\alpha|_{[0,t]}$. $J$ is open because $M$ is pseudo-special. Let $t_{0}:=\sup J$. For every $0\leq t<t_{0}$, there exists a unique continuous map $\beta_{t}:[0,t]\to X$ satisfying the above conditions. The map $\beta:[0,t_{0})\to X$ defined by $\beta(t):=\beta_{t}(t)$ is a continuous map by uniqueness of $\beta_{t}$. Let $(s_{m})_{m\in\mathbb{N}}$ be an increasing sequence in $0,t_{0})$ converging to $t_{0}$. By property $(\dagger)$, we may assume that $(\beta(s_{m}))_{m\in\mathbb{N}}$ is a convergent sequence by taking a subsequence. Put $c=\lim_{m\to\infty}\beta(s_{m})$. We have $\pi(c)=\alpha(t_{0})\in\pi(M)$. Therefore, we obtain $c\in\operatorname{cl}(X)\cap\pi^{-1}(\pi(M))$.
@@
We may assume that condition (i) is satisfied without loss of generality. By taking $u_{0}<t_{0}$ sufficiently close to $t_{0}$, we may assume that $\alpha([u_{0},t_{0}])$ is contained in $V^{\prime}$. By taking a subsequence of $(s_{m})_{m\in\mathbb{N}}$, we may assume that $\beta(s_{m})\in X\cap(Z^{\prime},C_{1})_{V^{\prime}}$ and $u_{0}<s_{m}<t_{0}$ for every $m\in\mathbb{N}$. By using the assumption that $M$ is a pseudo-special submanifold and condition (i), we can easily prove that $K:=\{t\in[s_{1},t_{0})\;|\;\beta(t)\in X\cap(Z^{\prime},C_{1})_{V^{\prime}}\}$ is closed and open in $s_{1},t_{0})$. We omit the details. Since $s_{1}\in K$, we have $K=[s_{1},t_{0})$. This means that $\beta([s_{1},t_{0}))\subseteq X\cap(Z^{\prime},C_{1})_{V^{\prime}}$. Since $\beta([s_{1},t_{0}))$ is connected, it is contained in a connected component of $X\cap(Z^{\prime},C_{1})_{V^{\prime}}$, say $D$. However, $D$ is a connected graph on $V^{\prime}$ with $Z^{\prime}<_{V^{\prime}}D$ by condition (i). We have $c\in\operatorname{cl}(D)$, but $c\in Z^{\prime}$ and $Z^{\prime}\cap\operatorname{cl}(D^{\prime})=\emptyset$, which is absurd.
@@
Assume for contradiction that $s_{0}:=\sup K<1$. Since $M$ is pseudo-special and uniqueness of $\widetilde{H}$, for every $0\leq t\leq 1$, there exist $\delta_{t}>s_{0}$ and $0\leq\varepsilon_{1,t}<t<\varepsilon_{2,t}\leq 1$ such that the restriction of $\widetilde{H}$ to $0,\delta_{t}]\times[\varepsilon_{1,t},\varepsilon_{2,t}]$ is continuous. Since $[0,1]$ is compact, there exists $\delta>s_{0}$ such that the restriction of $\widetilde{H}$ to $0,\delta]\times[0,1]$ is continuous, which is absurd. ∎
We show $\pi(X)=V$. Since $\pi|_{X}$ is a local homeomorphism, $\pi(X)$ is open in $V$. Assume for contradiction that $\pi(X)$ is not closed in $V$. We can find $a\in\operatorname{fr}(\pi(X))\cap V$. Let $(a_{m})_{m\in\mathbb{N}}$ be a convergent sequence in $\pi(X)$ converging to $a$. Take $b_{m}\in X$ with $\pi(b_{m})=a_{m}$. Since $M$ enjoys property $(\dagger)$, we may assume that $(b_{m})_{m\in\mathbb{N}}$ is convergent in $\mathbb{R}^{n}$ by taking a subsequence. Let $b:=\lim_{m\to\infty}b_{m}$. We have $\pi(b)=a\notin\pi(X)$ and $b\notin X$. This implies that $b\in\operatorname{fr}X\cap\pi^{-1}(\pi(M))$ and $\pi(b)\notin\pi(X)$. These contradict the either Claim 1 or ‘in particular’ part of Claim 2.
@@
We first reduce to the case where $k=1$. Put $A_{i}(0):=A_{i}$ and $A_{i}(1)=\mathbb{R}^{n}\setminus A_{i}$ for $0\leq i\leq k$. Let $\mathcal{S}$ be the set of sequences $\overline{s}$ of zeros and ones of length $k$ with $\overline{s}\neq(1,\ldots,1)$. Set $A_{\overline{s}}:=\bigcap_{i=1}^{k}A_{i}(s_{i})$ for $\overline{s}=(s_{1},\ldots,s_{k})\in\mathcal{S}$. By the hypothesis of reduction, for every $\overline{s}\in\mathcal{S}$, $A_{\overline{s}}$ is partitioned into finitely many multi-cells, say, $C_{\overline{s},1},\ldots,C_{\overline{s},l(\overline{s})}$. If $n=1$, $\{C_{\overline{s},i}\;|\;1\leq i\leq l(\overline{s}),\overline{s}\in\mathcal{S}\}$ is a desired multi-cell decomposition. If $n>1$, there exists a multi-cell decomposition partitioning $\{\pi(C_{\overline{s},i})\;|\;1\leq i\leq l(\overline{s}),\overline{s}\in\mathcal{S}\}$, say $D_{1},\ldots,D_{m}$. Let $\mathcal{D}$ be the collection of nonempty definable sets of the form $C_{\overline{s},i}\cap D_{j}$ for some $\overline{s}\in\mathcal{S}$, $1\leq i\leq l(\overline{s})$ and $1\leq j\leq m$. $\mathcal{D}$ is a desired multi-cell decomposition. Therefore, we may assume that $k=1$. We denote $A_{1}$ by $M$ for simplicity. We use this reduction process in this proof without notice.
We show this theorem by induction on $n$. Suppose $n=1$. Let $J=\operatorname{int}(M)$. Let $N=\operatorname{rank}(\operatorname{cl}(M\setminus J))$ and set $S_{i}:=\operatorname{iso}_{i}(\operatorname{cl}(M\setminus J))\cap(M\setminus J)$ for $0\leq i\leq N-1$. It is easy to show that $\{J,S_{0},\ldots,S_{N-1}\}$ is a multi-cell decomposition of $M$ after removing empty sets in the collection. We omit the details.
@@
Put $M[i]:=M\cap\pi^{-1}(\{x\in\pi(M)\;|\;\operatorname{rank}(\operatorname{cl}(M\cap\pi^{-1}(x)))=i\})$ for $1\leq i\leq N$. We may assume that $\operatorname{rank}(\operatorname{cl}(M\cap\pi^{-1}(x)))$ is independent of $x\in\pi(M)$, say $N$, by considering $M[i]$ instead of $M$. We define $P_{k}$ for $1\leq k\leq N-1$ as in clause (5) of Lemma 2.4. Put $P_{0}:=M$ for simplicity of notation. We prove the following claim which implies a multi-cell decomposition partitioning $M$.
Claim 1. There exists a multi-cell decomposition $\{C_{1},\ldots,C_{l}\}$ partitioning $\pi(M)$ such that, for $1\leq i\leq l$, $P_{i}\cap\pi^{-1}(C_{j})$ is a multi-cell for $0\leq i\leq N-1$ and clause (6) in Lemma 2.4 is satisfied for $M\cap\pi^{-1}(C_{i})$.
@@
Suppose $N>1$. Apply the induction hypothesis to $P_{1}$, then we may assume that $P_{1},\ldots,P_{N-1}$ are multi-cells and $(P_{i},P_{j})$ are extremely well-ordered pairs for $1\leq i<j\leq N-1$. Observe that $\pi(M)$ is a multi-cell because $\pi(M)=\pi(P_{1})$ and $P_{1}$ is a multi-cell. We want to reduce to the case where $(M,P_{i})$ is a well-ordered pair for every $1\leq i\leq N-1$. Put $T_{1i}:=\{x\in\pi(M)\;|\;\operatorname{Adh}^{+}(M,P_{i})\cap\pi^{-1}(x)\neq\operatorname{Adh}^{+}(M,P_{i},x)\}$ and $T_{2i}:=\{x\in\pi(M)\;|\;\operatorname{Adh}^{+}(M,P_{i})\cap\pi^{-1}(x)\neq\operatorname{Adh}^{+}(M,P_{i},x)\}$. Put $\bigcup_{i=1}^{N-1}T_{1i}\cup T_{2i}. TT$ is nowhere dense by Lemma 2.3.
@@
A multi-cell decomposition partitioning $\{V_{+\infty},V_{+\infty},V_{\mathbb{R}},W\}$ is a multi-cell decomposition partitioning $M$. Therefore, we have only to construct multi-cell decompositions partitioning each of $_{+\infty}, V+V_{+\infty}$, $V_{\mathbb{R}}$ and $W$. Observe that $V_{\mathbb{R}}=\pi(V_{\mathbb{R}})\times\mathbb{R}$. It is easy to show that $V_{\mathbb{R}}$ is a union of multi-cells. We omit the details. We next show that there exists a multi-cell decomposition partitioning $V_{+\infty}$. By considering a multi-cell decomposition of $\pi(V_{+\infty})$, we may reduce to the case where $\pi(V_{+\infty})$ is a multi-cell. Consider the definable function $f:\pi(V_{+\infty})\to\mathbb{R}$ given by $f(x)=\inf\{t\in\mathbb{R}\;|\;(x,t)\in V_{+\infty}\}$. Let $D$ be the set of points at which $f$ is discontinuous. By [5, 8.3 Almost continuity], we have $\dim D<\dim\pi(V_{+\infty})$. We can reduce to the case in which $f$ is continuous by induction on $\dim\pi(V_{+\infty})$. $V_{+\infty}$ is a multi-cell because $\pi(V_{+\infty})$ is a multi-cell and $V_{+\infty}=\{(x,t)\in\pi(V_{+\infty})\times\mathbb{R}\;|\;t>f(x)\}$. We can prove that there exists a multi-cell decomposition partitioning $V_{-\infty}$ in the same manner.
@@
Let $Q:=\{(x,y)\in\pi(W)\times\mathbb{R}\;|\;\exists z\ (x,y,z)\in Z\}$. Observe that $\pi(W)=\pi(Q)$ and $\dim Q\cap\pi^{-1}(x)=0$ for every $x\in\pi(Q)$. By Claim 1, there exists a multi-cell decomposition $\{Q_{1},\ldots,Q_{l}\}$ of $Q$. Observe that, for every $(x,y)\in Q$, there exists a unique $z\in\mathbb{R}$ with $(x,y,z)\in Z$. Let $g:Q\to\mathbb{R}$ be the definable function sending $(x,y)\in Q$ to the unique point $z\in\mathbb{R}$ with $(x,y,z)\in Z$. Put $g_{i}:=g|_{Q_{i}}$ for $1\leq i\leq l$. By [5, 8.3 Almost continuity], the set $G_{i}$ of points at which $g_{i}$ is discontinuous is nowhere dense in $Q_{i}$. For every $\in\operatorname{Con}(Q_{i}), CC$ is the graph of a continuous function defined on a connected component of $\pi(Q_{i})$. Therefore, $\pi(G_{i}\cap C)$ is nowhere dense in $\pi(Q_{i})$ for every $C\in\operatorname{Con}(Q_{i})$. Since $\operatorname{Con}(Q_{i})$ is a countable set, $\pi(G_{i})=\bigcup_{C\in\operatorname{Con}(Q_{i})}\pi(G_{i}\cap C)$ has an empty interior in $\pi(Q_{i})$ by the Baire category theorem. Since $\pi(G_{i})$ is definable in the d-minimal structure $\mathfrak{R}$, $\dim\pi(G_{i})<\dim\pi(Q_{i})$. By induction on $\dim(\pi(Q_{i}))$, using the induction hypothesis, there exists a multi-cell decomposition $D_{1},\ldots,D_{m}$ of $\pi(Q)$ such that $Q\cap\pi^{-1}(D_{i})$ is a multi-cell and the restriction of $g$ to $Q\cap\pi^{-1}(D_{i})$ is continuous. Since $Q\cap\pi^{-1}(D_{i})$ is a multi-cell, every connected component of $W\cap\pi^{-1}(D_{i})$ is of the form $\{(x,y)\in D_{i}\times\mathbb{R}\;|\;f(x)<y<g(x,f(x))\}$, where $f$ is a continuous function defined on $D_{i}$. This means that $W\cap\pi^{-1}(D_{i})$ is a multi-cell. ∎
@@
Let $\mathfrak{R}$ be a d-minimal expansion (,<,+,)($\mathbb{R}$$\ldots$) of the ordered group of reals which satisfies one of the following conditions: